OpenStudy (anonymous):

simpify 2log5 12-log5 8-2log5 3

OpenStudy (anonymous):

first we can make: 2 log5 12 = log5 12^2 = log5 144, similarly 2 log5 3 = log5 9 so we have: log5 144 - log5 8 = log5 (144/8) = log5 18 log5 18 - log5 9 = log5 (18/9) = log5 2 the answer is: -------------------------------------------- | 2log5 12-log5 8-2log5 3 = log5 2 | --------------------------------------------